### v2

# Navigating the Latent Space Dynamics of Neural Models

Abstract
Neural Network (NN) transform high-dimensional data into compact, structured representations, often modeled as elements of a lower dimensional latent space. In this paper, we present an alternative interpretation of neural models as dynamical systems acting on the latent manifold. Specifically, we show that AutoEncoder (AE) models implicitly define a latent vector field on the manifold, derived by iteratively applying the encoding-decoding map, without any additional training. We observe that standard training procedures introduce inductive biases that lead to the emergence of attractor points within this vector field. Drawing on this insight, we propose to leverage the vector field as a representation for the network, providing a novel tool to analyze the properties of the model and the data. This representation enables to: $(i)$ analyze the generalization and memorization regimes of neural models, even throughout training; $(ii)$ extract prior knowledge encoded in the network’s parameters from the attractors, without requiring any input data; $(iii)$ identify out-of-distribution samples from their trajectories in the vector field. We further validate our approach on vision foundation models, showcasing the applicability and effectiveness of our method in real-world scenarios.
1 Introduction
Figure 1: The latent dynamics of AEs Latent vector fields with $\mathbf{z}_{0}\sim\mathcal{U}[-8,8]$ of 2D AEs trained on the MNIST dataset with different initializations. Colors denote the norm of the vectors: the boundary denotes the shape of the latent manifold identified with the support of the encoder, in white we observe regions of the latent space where the vector field vanishes, flowing toward attractors, which correspond to regions of high probability density of the data.
Neural models are effective function approximators that have been demonstrated to be able to solve complex tasks in countless domains [23, 41]. These models typically learn to transform high-dimensional data into compact, structured representations [4], often modeled as elements of a lower dimensional latent space. In this work, we propose a novel alternative representation for neural models by showing that, for a given AE architecture, there exists an associated latent vector field (see Figure 1), which characterizes the behavior of the network at the representation level. The vector field is obtained by simply iterating the encoding decoding map in the latent space, without any additional training procedures. Previous work [42, 22] has demonstrated that in the extreme overparametrized regime and under strong assumptions, AEs can effectively memorize training data. In this work, we propose to consider the more general case in which memorization corresponds to one extremum, showing evidence that the latent vector field induced by general AEs can be reconnected to properties of the network, such as generalization, memorization, and out-of-distribution detection. Our contributions can be summarized as follows:
  • We show that every AE architecture induces a latent vector field, implicitly. We demonstrate that this representation retains properties of the networks and the data, throughout its trajectories and its fixed points.
  • We show that most mappings learned by NNs are contractive: this enforces the presence of fixed points and attractors points in the latent vector field.
  • We empirically show how attractors characterize the memorization and generalization regime in NNs and their interplay in the training process.
  • We experimentally demonstrate the information stored in vision foundation models weights can be retrieved in a data-free way, by probing models with noise and retrieve attractors.
  • Finally, we show that paths in the latent vector fields are informative to characterize the learned distribution and to detect distribution shifts.
2 Method
Notation and background
In the following sections we will consider neural models $F$ as a composition of encoder decoder modules $F_{\Theta}=D_{\theta_{2}}\circ E_{\theta_{1}}$ parametrized by $\Theta=[\theta_{1},\theta_{2}]$. The encoder maps samples $\mathbf{x}\sim p(\mathbf{x})$ supported on $\mathcal{X}\subset\mathbb{R}^{m}$ to a typically lower dimensional space $\mathcal{Z}\subset\mathbb{R}^{k}$ and the network $F$ is trained for reconstruction to minimize the mean squared error $\mathcal{L}_{MSE}$:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in X}\|\mathbf{x}-F_{\Theta}(\mathbf{x})\|_{2}^{2}+\lambda\mathcal{R}(\Theta)$ (1)
where $\mathcal{R}$ is either an explicit (e.g., weight decay) or implicit regularization constraint on the parameters $\Theta$ to favor contractive solutions of $F$. We posit that minimizers of the above objective act as minimizers of the spectral norm of the Jacobian $\rho(J_{F}(x))$ for $x\sim p(x)$. Some examples of regularizer include weight decay, where $\mathcal{R}=\|\Theta\|_{2}^{F}$ or the use of some augmentations distribution $p(T)$, for which the objectives becomes:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in X}\|\mathbf{x}-F(\mathbf{x})\|_{2}^{2}+\sum_{T\in p(T)}\|\mathbf{x}-F(T\mathbf{x})\|_{2}^{2}$ (2)
examples of $p(T)$ include additive gaussian noise in denoising AEs [52] or masking, in masked AEs [15]. The bottleneck dimension $k=dim(\mathcal{Z})$ is another example of how to enforce a contractive behavior in the network. More precisely, the bottleneck dimension acts as a hard constraint on the rank of the Jacobian of the encoder $rank(J_{E})<=k$ and therefore of the entire map $F$, since $J_{F}(\cdot)=J_{E}(\cdot)\odot J_{D}(\cdot)$. In Table C in the Appendix, we list many AE variants including denoising AEs (DAEs) [52], sparse AEs (SAEs) [37], variational AEs (VAEs)  [26] and other variants [44, 2, 12] and show how their objectives enforce local contractive solutions around training points.
Given an AE model $F$ possibly pretrained, denoting with $f(\mathbf{z})=E\circ D(\mathbf{z})$, we are interested in analyzing the iterations of the map $f(\ldots f(f(\mathbf{z})))$, which can be modeled as a differential equation.
2.1 The Latent Dynamics of Neural Models
For a sample $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{k}$ we analyze the effect of applying iteratively $f$ on $\mathbf{\mathbf{z}}$, i.e. $f\circ f\circ\ldots f(\mathbf{z})$. Parametrizing the number of iteration by $t$, this lead to the following discrete ordinary differential equation (ODE):
$\begin{cases}&\mathbf{z}_{t+1}=f(\mathbf{z})\\ &\mathbf{z}_{0}=\mathbf{z}\end{cases}$ (3)
which discretizes the following continuous differential equation:
$\frac{\partial\mathbf{z}}{\partial t}=f(\mathbf{z})-\mathbf{z}$ (4)
In Eq 3, one can associate the map $f$ as the pushforward of a latent vector field $V:\mathbb{R}^{k}\mapsto\mathbb{R}^{k}$ tracing non linear trajectories in the latent space (see Fig 1). A first interesting question is whether this ODE has solutions and if they are unique. Through the Banach fixed point theorem, it can be shown that the ODE has a unique solution iff the map $f$ is Lipschitz continuous.
Definition 1.
A function $f:\mathcal{Z}\mapsto\mathcal{Z}$ is denoted Lipschitz continuous if there exist a constant $C$ s.t. for every pair of points $\mathbf{x_{1}},\mathbf{x_{2}}$:
$d(\mathbf{z_{1}},\mathbf{z_{2}})_{\mathcal{X}}\leq C\ d(f(\mathbf{z_{1}}),f(\mathbf{z_{2}}))_{\mathcal{X}}$ (5)
when $C<1$ $f$ is called contractive.
For any contractive map, there will exist fixed point solutions of Eq 3, i.e., iteratively applying the map $f$ will converge to a unique constant solution via solving $\mathbf{z}^{*}=f(\mathbf{z}^{*})$. The fixed point $\mathbf{x}$ solutions correspond to singular points that attract and summarize the dynamics of the system.
Definition 2.
A fixed point $\mathbf{x}=f(\mathbf{x})$ is an attractor of a differentiable map $f$ if all eigenvalues of the Jacobian $J$ of $f$ at $\mathbf{x}$ are strictly less than one in absolute value.
When $f$ is nonlinear, the ODE in Eq. 3 can have multiple solutions depending on the initial conditions $\mathbf{z_{0}}$. The previous definition allows for a definition of Lipschitz continuity which is inherently local:
Definition 3.
[…]
The set of initial conditions $Z_{0}$ leading to the same attractor $\mathbf{z}^{*}$, is denoted as basin of attraction.
Why maps learned by NNs are contractive in practice? We argue that most maps learned by neural AEs or variants thereof, promote (locally) contractive autoencoding mappings. In other words, many explicit or implicit regularization strategies and training design choices constrain the Jacobian of the autoencoding map to have small eigenvalues near training examples. We provide below what are the inductive biases that promote locally contractive mappings:
  • Initialization Standard initialization schemes [29, 14, 16] aim at preserving the variance of the activations at initialization to avoid gradient vanishing and exploding problems during training. Theoretical arguments in support assume i.i.d. input samples and weights, and are generally tied to specific activation functions (e.g., ReLU in [16]). We note that in practice, this condition is not likely to hold, as training data samples are often correlated and many employed architectural elements (for example, residual connections [17]) break the independence assumptions of the layer weights. Networks result in a bias to be either expansive or contractive towards a single point, with a preference for the latter case [39]. We show empirically that this is indeed the case in Figure 6 in the Appendix, across several vision backbone models. In Figure 3, we show an example of a 2D latent vector field which is globally contractive at initialization (Figure 3(a), left plot) towards a single attractor at the origin. This is the case even in higher dimensions (see the number of attractors at the first epoch in Figure 3(c)).
  • Explicit regularization Widely employed regularization techniques such as weight decay [7], implicitly constraints the learned map to be (locally) contractive. This can be seen easily in the case of linear mappings, where penalizing the norm of the matrix corresponds to penalizing the first-order derivatives of the map, resulting in the optimal way to make the map contractive. Although this is not guaranteed to be optimal for nonlinear networks, weight decay still enforces a contraction mapping. The regularizer is incorporated in state-of-the-art optimizers [32], ubiquitous in state-of-the-art foundation models. Other forms of parametric constraints include the choice of bottleneck dimension, which acts as a hard constraint on the rank of $J_{F}(\mathbf{z})$, and soft variants thereof, such as the Kullback-Leibler divergence term in VAEs, or sparsity regularization in SAEs.
  • Implicit regularization As previously discussed, the choice of augmentation distribution implicitly selects the notion of neighborhood of the training samples (for example, Gaussian radius in DAEs), and the perturbation direction that will be penalized in $J_{f}(\mathbf{x})$. This form of regularization is nonparametric and inherently local as opposed to the explicit regularization techniques explained in the above paragraph.
3 Theoretical remarks
In this section, we are interested in analyzing properties of trajectories within the latent vector field, as well as its fixed points and attractors, when they exist.
3.1 Trajectories of the latent vector field
In [2, 34, 45], it was shown that, under optimal conditions, the residual in denoising AEs trained with noise variance $\sigma^{2}$ approximates the score function, i.e., the gradient of the log empirical density of the data, as $\sigma\rightarrow 0$. Interestingly, [2] shows a connection between denoising and contractive AEs, where the norm of the Jacobian of the input-output mapping is penalized. The more general result we want to stress here is that whenever the auto-encoding map is locally contractive w.r.t. a given notion of neighborood (e.g. gaussian radius, masking augmentations) and approximates well enough the input distribution $p(\mathbf{x})$, then the associated latent vector field pushes samples in the direction of the derivative of the score of the approximated prior in the latent space. Informally, this means that the vector field will non-linearly project samples on the data manifold onto regions of high probability density.
Theorem 1.
[…]
In general, attractors cannot be computed via solving the fixed point equation with first-order methods, unless the initial condition is arbitrary close to the attractors, as the following result shows:
Proposition 3.1.
Iterations of the map $f$ in Equation 3 correspond to performing gradient descent in the $\frac{\partial L(\mathbf{z})}{{\partial\mathbf{z}}}$ iff $f$ is an isometry (i.e. its eigenvalues are concentrated near 1) or in the proximity of attractor points, when the Jacobian of $f$ vanishes.
This result highlights the importance of the higher order terms involved in iterating $f$, which trace a highly nonlinear trajectory in latent space, escaping local minima solutions given by unstable and spurious fixed points.
3.2 Characterizing the attractors of the latent vector field
In this section we aim to characterize what the fixed point solution of Eq 3 represents. We start by considering two base cases:
Linear maps In the linear case, the encoding map $E$ and decoding $D$ are parametrized by matrices $\mathbf{W}_{1}\in\mathbb{R}^{N\times k}$ and $\mathbf{W}_{2}\in\mathbb{R}^{k\times N}$ and the only fixed point of the map corresponds to the origin if the network is bias free or to a shift of it. The rate of convergence of the iteration in Eq 3 is established by the spectrum of $\mathbf{W}_{2}^{T}\mathbf{W}_{1}$, and the iteration is equivalent to shrinking the input in the direction of the eigenvectors with associated eigenvalue $\lambda<1$. In case the eigenvectors of the trained AEs are aligned with the optimal solution given by the Principal Component Analysis (PCA) decomposition of the data $\mathbf{X}=\mathbf{\Phi}\mathbf{\Lambda}\mathbf{\Phi}^{*}$, then the latent vector field vanished as the mapping is isometric and no contraction occurs.
Homogeneous maps In the case of homogeneous NNs, the network satisfies $F(c\mathbf{x})=c^{\alpha}F(\mathbf{x})$ with $c\in\mathbb{R}$ for some $\alpha$. For example, this holds for ReLU networks without biases with $\alpha=1$, which learn a piecewise linear mapping. The input-output mapping can be rewritten as:
$F(\mathbf{x})=J_{F}(\mathbf{x})\mathbf{x}$ (6)
A similar observation was made in for denoising networks in [24], we remark here its generality.
This equality implies that we can rewrite Equation 3 as :
$\mathbf{z}_{t+1}=J_{f}(\mathbf{z}_{t})\mathbf{z}_{t}=\sum_{i}\lambda_{i}\mathbf{\phi}_{i}\mathbf{z}_{t}$ (7)
Where $\sum_{i}\lambda_{i}\mathbf{\phi}_{i}$ is the eigendecomposition of $J_{f}(\mathbf{z})$. Since in the proximity of an attractor $max\lambda(J_{F}(\mathbf{z}^{*})<=1$, the iterations shrink directions corresponding to the eigenvectors of the Jacobian by the corresponding eigenvalue. This allows us to derive the following result on the speed of convergence.
Proposition 3.2.
The error $e_{t}=\|\mathbf{z}_{t}-\mathbf{z}^{*}\|$ converge exponentially to an $\epsilon$ depending on the spectral norm $\rho(J_{f}(\mathbf{z}^{*}))$, according to the formula $\frac{\log(\frac{\epsilon}{\|\mathbf{e}_{0}\|})}{\log(\|\rho(J_{f}(\mathbf{z}^{*}))\|)}$ that provides an estimate for the number of iterations $T$ to converge to the attractor.
In Figure 11 in the Appendix, we measure how well the convergence formula predicts the measured number of iterations to converge to an attractor, on a fully non linear AEs with biases, where the assumptions of the current section are less likely to hold. The error in the estimate will be higher, when the initial condition $\mathbf{z}_{0}$ is far from the attractor, as higher order terms dominate the dynamics, and the first order Taylor approximation accumulates more error.
Extreme overparametrization case When the capacity of a network exceeds the number of training examples by far, AEs enter in a overfitting regime which leads to data memorization [56, 42, 22, 24]. The network in this case learns a constant function or an approximation thereof, which can be retrieved through the iterations in Eq:3.
Between memorization and generalization
[…]
(a) Attractors
Figure 2: Memorization vs Generalization: Attractors memorize the training data as a function of the rank of $J_{f}(\mathbf{z})$ by adjusting the bottleneck dimension $k$ (Left); which is inversely proportional to the amount of generalization attained by the model (center); On the right we show example of attractors transition from a strong memorization model (First row), to good generalization (Last row).
Our goal is to characterize the properties of the latent vector field and the learned attractors in the general setting. We argue that different models can perfectly fit training points reaching similar low loss, but interpolate differently outside the training support, depending on the regularization strength and the amount of overparametrization. In this context, which information do $\mathbf{z}^{*}$ store in general? We argue that neural models fall in a spectrum between memorization and generalization, depending on the strength of the regularization term in 1, and we hypothesize that attractors fully characterize where in the spectrum a model is. We consider attractors as representations summarizing the information stored in the weights of the network. This interpretation is in line with the convergences to modes of Theorem 3.1 and can be easily seen in the case of memorization. Specifically, we assume that attractors form a dictionary of $N$ signals $\mathcal{D}^{*}=\{\mathbf{z}_{1}^{*},...,\mathbf{z}_{P}^{*}\}$ from a set of initial conditions sampled from an input distribution $P_{0}$. The dictionary corresponds to, possibly nonlinear, interpolations of the training data X, therefore $\mathcal{D}^{*}=\Phi(X)$. In the memorization regime, $\Phi$ corresponds to the standard basis up to permutations, and the network approximates a constant function which can be retrieved through the iterations of Equation 3. Transitioning to a generalization regime $\Phi$ can become arbitrarily complex. We demonstrate the interplay between memorization and generalization empirically in the experimental sections below. In Figure 2, we show how a model transitions from memorization to generalization according to the strength of the regularization on $J_{f}(\mathbf{z}^{*})$, by measuring how many training examples are stored as attractors. We remark that this memorization corresponds to an over regularized regime, as opposed to the overfitting regime of extremely overparametrized networks.
Setting We trained 30 convolutional AEs models on the CIFAR, MNIST, and FashionMNIST datasets, varying the bottleneck dimension $k$ from 2 to 512. This acts as a hard regularizer on $J_{f}(\mathbf{x}^{*})$, by constraining its rank to be $<=k$. We compute attractors $\mathbf{Z}^{*}$ from elements of the training set $\mathcal{X}_{train}$ by iterating $\mathbf{z}$ till convergence. We measure the degree of memorization by defining a memorization coefficient, which is given by the cosine similarity of each decoded attractor $\mathbf{x}^{*}=D(\mathbf{z}^{*})$ to its closest point in the training set, $mem(\mathbf{z}^{*})=\min_{x\in\mathcal{X}_{train}}cos(D(\mathbf{z}^{*}),\mathbf{x})$ and we report the mean and standard deviation over attractors. We measure generalization by simply reporting the error on the test set. In Figure 12 in the Appendix, we also report the rank explaining $90\%$ of the variance of the matrix of decoded attractors $\mathbf{X}^{*}$, showing that attractors corresponding to good generalization models span more directions in the input space. For additional information on the model, hyperparameters, and settings, we refer to the Appendix.
Analysis of results In Figure 2 on the left, we observe that the more regularized networks ($k$ from 2 to 16) tend to memorize data, trading off generalization performance (middle plot). We remark that this kind of memorization is not due to overfitting, as was shown in [24] for diffusion models, but it happens in the underfit regime, due to the strong regularization constraint on $J_{f}(\mathbf{x})$. We show in the Appendix results in the overfitting regime, by training the models with different sample sizes and observing a similar pattern.
Takeaway: Attractors capture the interplay between generalization and memorization of neural models, which corresponds to the trade-off between the reconstruction performance and regularization term of the AEs model.
[…]
(a) Evolution of a 2D Latent Vector field over training iterations
(b) Memorization to generalization
(c) Number of attractors
(d) Similarity $\mathbf{Z}_{noise}^{*}$ and $\mathbf{Z}_{train}^{*}$
Figure 3: The dynamics of the latent vector field (a) The 2D vector field expands from a single attractor, eventually stabilizing and over-fitting because of capacity limits. Bottom: Evolution of larger capacity AEs ($k=128)$ across training. (b) Throughout training, the network first memorizes the data with a high memorization coefficient (in blue) and then generalizes, achieving a low test error (red). (c): Evolution of attractor count for training (blue), test (red), and noise (yellow) samples; (d) Attractors computed from training and from gaussian noise converge during training (green), while the separability of the trajectories (measured as FPR95, the lower the better) increase (purple).
In Figure 3 we show a similar transition from a memorization regime to generalization across training. Setting We monitor the latent vector field and attractors statistics across the training dynamics of a convolutional AEs trained on the MNIST with bottleneck dimension $k=128$.
Analysis of results In Figure 3(b) we show the transition from memorization, occurring at the first epochs of training, to generalization, by plotting the memorization coefficient and the test error across training, observing a trade off between the two. In the central plot, we plot the fraction of distinct attractors, computed from 5000 random elements respectively from the $P_{train},P_{test}$ and $\mathcal{N}(0,I)$, where we consider two attractors equal if $cos(\mathbf{z}_{1}^{*},\mathbf{z}_{2}^{*})>0.99$. We observe that the number of attractors grows across epochs starting from a single attractor at initialization, as we can also observe in the 2d case in Figure 3(a), with the training and test number of attractors stabilizing according to the test loss and the noise ones converging later. On the right, we plot the similarity of attractors computed from the training and the noise distributions, measured by Chamfer symmetric similarity, observing that it increases across training, as the two distributions of attractors match. Despite being able to compute similar attractors starting from noise or training points, the trajectories in latent vector to attractors are very different, according to the source distribution considered: we show this by reporting FPR95 scores for separating trajectories from noise distribution to train distribution, based on the following measure $score(\mathbf{z})=\frac{1}{N}\sum_{z_{i}\in\pi(\mathbf{z})}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$ where $\pi(\mathbf{z})=[\mathbf{z}_{0},...,\mathbf{z}_{N}]$ is the trajectory. FPR drastically decreases with training, demonstrating that the networks achieves separability of the two distributions of trajectories.
Takeaway: As the latent vector field evolves during training, the model transitions from memorization to generalization, forming similar attractors from different input distributions, while retaining information of the source distribution in the latent trajectories.
4 Experiments on vision foundation models
In this section we demonstrate the existence and applicability of the latent dynamics on vision foundation models, including AE backbone of the Stable Diffusion model[46] variants and vision transformer masked AEs [15], showing; (i) how information stored in the weights of a pretrained model can be recovered by computing the attractors from noise (ii) how trajectories in the latent dynamics are informative to characterize the learned distribution and detect distribution shifts.
4.1 Data free weight probing
[…]
(a) Reconstructions with $5\%$ atoms. First row: Input; second row: Orthogonal basis; last row: Attractors
Figure 4: Data-free weight information probing of Stable Diffusion model: We plot the error (MSE) vs sparsity (number of atoms) used to reconstruct samples from diverse dataset respectively from (i) an orthonormal random basis of the latent space (blue); (ii) attractors computed from gaussian noise (red), showing that attractors consistently reconstructs samples better on all datasets. (Right) Reconstructions using $5\%$ of the atoms on ImageNet
Setting. In the following experiment, our goal is to analyze to what degree we can recover information stored in the weights of a neural model from attractors computed from Gaussian noise, without relying on any input data. We focus on the analysis of the AE of stable diffusion [46] pretrained on the large scale Laion2B dataset [47]. We sample input $\mathbf{Z}_{n}\sim\mathcal{N}(\mathbf{0},I)$ and compute the attractors corresponding to the solution of the equation $f(\mathbf{Z}_{n}^{*})=\mathbf{Z}_{n}^{*}$. Specifically, we sample $N=4096$ points which correspond to the latent dimensionality of the model $k$. We evaluate the reconstruction performance on 6 different and diverse datasets Laion2B [47],Imagenet [9], EuroSAT [18], CIFAR100 [27],PatchCamelyon [30], Places365 [31], showcasing highly different distributions of images from general domains, to medical and satellite data. We evaluate on 500 test samples from each dataset, sampled randomly, using Orthogonal Matching Pursuit (OMP) [33] to reconstruct the sample from the attractors in the latent space, as a function of the sparsity imposed by the number of atoms selected. We compare with the reconstructions from the initial orthogonal basis $Z_{0}^{*}$, which, when using all samples, span completely $\mathcal{Z}$. The reconstruction performance is measured in the output space, decoding the reconstructed latents via $D$.
[…]
Takeaway. Attractors of foundation models computed from noise can serve as a dictionary of signals to represent diverse datasets, demonstrating that it is possible to probe the information stored in the weights of foundation models in a black box way, without requiring any input data.
4.2 Latent trajectories characterize the learned distribution
In the following experiment, we test the hypothesis that trajectories are informative on the distribution learned by the model, testing how much Theorem 3.1 holds in real cases. In particular, we want to assess how much the paths in the vector field can detect distribution shifts in the data. For a sample to be classified as out-of-distribution (OOD) we identify two relevant questions to investigate: (i) does the sample iteration converge to one of the attractors of the training data, i.e., does it share the same basin of attraction? (ii) If so, how fast does it converge to the attractor? An OOD sample can, in fact, lie inside one of the attractor basins shared with training data. In order to assess this, we propose to track the distribution of distances to attractors point along the encoding-decoding iteration of a test sample. In the former case, we expect that an OOD sample converges faster, while in the latter case, the distance term to the attractors will dominate the score.
[…]
Analysis of results. In Table we report OOD detection performance by using the following score function: for a test sample $\mathbf{z}_{test}$ we compute its trajectory $\pi_{\mathbf{z}_{test}}=[\mathbf{z}_{0},...,\mathbf{z}_{N}^{*}]$ towards attractors and compute the distance of $\pi_{\mathbf{Z}_{test}}$ w.r.t. to the set of training attractors $\mathbf{Z}_{train}^{*}$, i.e. $score(\mathbf{z})=d(\pi_{\mathbf{z}_{test}},\mathbf{Z}_{train}^{*})$. As a distance function, we employ Euclidean distance, and we aggregate the score computing the mean distance over the trajectory $score(\mathbf{z})=\frac{1}{N}\sum_{i}d(\mathbf{z}_{i},\mathbf{Z}_{train}^{*})$. We compare with a $K$-Nearest neighbor baseline, where the score for a test sample is obtained by taking the mean distance over the $K$-NN on the training dataset, where $K=2000$ in the experiments. The score proposed demonstrates how informative the latent vector field is on the training distribution. In Figure 5 on the right, we show histograms of scores for the distance to attractors and the KNN, showing again that the former method is able to tell apart in-distribution and out-of-distribution data correctly.
Takeaway. Trajectories in the latent vector field characterize the source distribution and are informative to detect distribution shifts.
SUN397 Places365 Texture iNaturalist
[…]
d(Attractors) 29.60 91.20 29.95 90.99 25.85 92.63 29.85 91.29
KNN 100.00 42.59 100.00 32.36 34.50 89.41 86.35 68.60
(a) FPR and AUROC scores
Figure 5: Trajectories in the latent vector field characterize distribution shifts We measure out-of-distribution detection performance on ViTMAE: On the left we report scores for 4 different datasets, highly outperforming the KNN baseline. On the right histograms of scores the INaturalist dataset demonstrating much better separability between in-distribution and out-of-distribution.
5 Related Work
Memorization and generalization in NNs NNs exhibit a rich spectrum of behaviors between memorization and generalization, depending on model capacity, regularization, and data availability [3, 57, 40]. In the case of extreme overparametrization, namely networks trained on few data points, it has been showed experimentally in [42, 56] and theoretically for sigmoid shallow AEs in [22] that AEs can memorize examples and implement associative memory mechanisms. [2, 51], A similar phenomenon has been observed in diffusion models in [48, 8] and analyzed in [24] Similarly, non gradient-based approaches such as Hopfield networks and their modern variants [19, 43] extend classical attractor dynamics to neural system that interpolate between memory-based and generalizing regimes. In our work, we show that AEs fall in general in the spectrum between memorization and generalization, depending on inductive biases that enforce contraction.
Contractive neural models: Different approaches have been proposed to regularize NNs in order to make them smoother and more robust to input perturbations and less prone to overfitting. Many of these regularization techniques either implicitly or explicitly promote contractive mappings in AEs: for example sparse AEs [37, 12], their denoising variant [52], and contractive AEs [44, 2] losses enforce learned maps to be contractive through the loss. Regularization strategies such as weight decay [28] favor as well contractive solutions for neural models, and they are incorporated in standardized optimizers such as AdamW [32]. All these approaches enforce either directly or indirectly the existence of fixed points and attractors in the proposed latent vector field representation.
Nonlinear operators spectral analysis In the context of image processing and 3D graphics, previous work has inspected generalization of spectral decompositions to nonlinear operators [5, 13, 10], focusing on one homogeneous operator (e.g., p-Laplacians). In our method, the fixed points computation of the vector field can be interpreted as the decomposition of the NN into a dictionary of signals
6 Conclusions and Discussion
In this manuscript we proposed to represent neural AEs as vectors fields, implicitly defined by iterating the decoding encoding map in the latent space. We showed that attractors in the latent vectors fields exists in practice due to inductive biases in the training regime which enforce local contractive solutions and they retain important properties of the model and the data, linking to memorization and generalization properties of the model. We show that the information stored in the weights can be retrieved without using any input data in vision foundation models and paths in the vector field are informative to characterize the learn distribution and detect distribution shifts.
Limitations and future works
Generalizing to arbitrary models Eq. 3 cannot be directly generalized to model trained with discriminative objectives such as a deep classifiers, or self supervised models, as the network is not invertible. However we note that point space the vector field is still defined in the output space, and can be simply by computing the residual $F(\mathbf{x})-y$. and in the neighborhood of an attractor, the relation in proposition 1 can still hold for different objectives.An alternative intriguing idea is the one to train a surrogate AE model in the latent space of the model of interest, which would be agnostic from the pretraining objective. Sparse AEs for mechanistic interpretability of large language models (LLMs) [12] fit in this category. Analyzing the associated latent vector field can shed light on features learned by SAEs and biases stored in their weights.
Learning dynamics: characterizing how attractors forms during training, under which conditions noise attractors converge to the training attractors, holds promise to use the proposed representation to study the learning dynamics of neural models to inspect finetuning of AE modules, such as low-rank adapters [20] and double descent [36].
Alignment of latent vector field: Finally, following recent findings in representation alignment [35, 11, 21] inspecting how latent vector fields of networks trained are related is an open question for future works, to possibly use this representations to compare different neural models.
Acknowledgments and Disclosure of Funding
This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/COE12. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. MF is supported as well by the MSCA IST-Bridge fellowship which has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101034413. M.F. thanks C.Domine, V.Maiorca, I.Cannistraci, R.Cadei, B.Demirel, S.Vedula for insightful discussions.
Appendix A Proofs and derivation
A.1 Theorem 3.1
We report here a detailed formulation of Theorem 3.1 alongside its proof.
Assumption A.1 (Latent marginal).
Let $p_{\mathrm{data}}(\mathbf{x})$ be the data distribution and $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ an encoder mapping inputs $\mathbf{x}\in\mathbb{R}^{n}$ to latent codes $\mathbf{z}\in\mathbb{R}^{d}$. The latent marginal density is defined as:
[…]
We assume that $q(\mathbf{z})$ is continuously differentiable, and that its gradient $\nabla\log q(\mathbf{z})$ and Hessian $\nabla^{2}\log q(\mathbf{z})$ are well-defined and continuous on an open domain containing $\Omega\subseteq\mathbb{R}^{d}$.
Assumption A.2 (Fixed-point manifold).
Let $E:\mathbb{R}^{n}\to\mathbb{R}^{d}$ be an encoder and $D:\mathbb{R}^{d}\to\mathbb{R}^{n}$ a decoder, both continuously differentiable. Define the composite map
$f(\mathbf{z})=E(D(\mathbf{z}))\in\mathbb{R}^{d}.$
We define the fixed-point manifold of $f$ as
=$\mathcal{M}=\left\{\mathbf{z}\in\mathbb{R}^{d}\;:\;f(\mathbf{z})=\mathbf{z}\right\}.$
Assumption A.3 (Local contraction).
There exists an open, convex set $\Omega\subseteq\mathbb{R}^{d}$, with $\mathcal{M}\subseteq\Omega$, and a constant $10<L<$ 1, such that:
$\sup_{\mathbf{z}\in\Omega}\left\|J_{f}(\mathbf{z})\right\|_{\mathrm{op}}\leq L,$
where $J_{f}(\mathbf{z})\in\mathbb{R}^{d\times d}$ denotes the Jacobian of $f$ at $\mathbf{z}$, and $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm (i.e., largest singular value). Therefore, $f$ is a contraction mapping on $\Omega$.
Assumption A.4 (Training optimality with Jacobian regularization).
Let $f(\mathbf{z})=E(D(\mathbf{z}))$ be as above, with $E,D$ trained to minimize the objective:
$\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}}\left[\|D(E(\mathbf{x}))-\mathbf{x}\|^{2}+\lambda\|J_{f}(E(\mathbf{x}))\|_{F}^{2}\right]\quad\text{for }\lambda>0.$
Assume the encoder is deterministic, i.e., $q_{\phi}(\mathbf{z}|\mathbf{x})=\delta(\mathbf{z}-E(\mathbf{x}))$, so the latent marginal satisfies:
[…]
Then $q(\mathbf{z})$ is supported and concentrated on the fixed-point manifold $\mathcal{M}$, and $f$ is locally contractive around $\mathcal{M}$.
Lemma A.5 (Directional ascent of the residual field).
Under the above assumptions, for every $\mathbf{z}\in\Omega\setminus\mathcal{M}$, define the residual field:
$\mathbf{v}(\mathbf{z}):=f(\mathbf{z})-\mathbf{z}.$
Then the directional derivative of $\log q$ in direction $\mathbf{v}(\mathbf{z})$ is strictly positive:
$\left\langle\nabla\log q(\mathbf{z}),\mathbf{v}(\mathbf{z})\right\rangle>0.$
Proof.
Let $\mathbf{z}\in\Omega\setminus\mathcal{M}$ be arbitrary. Then $f(\mathbf{z})\neq\mathbf{z}$, so $\mathbf{v}(\mathbf{z})=f(\mathbf{z})-\mathbf{z}\neq\mathbf{0}$. By contractiveness and training optimality, $f(\mathbf{z})$ is closer to the fixed-point set $\mathcal{M}$ than $\mathbf{z}$ is. Since $q(\mathbf{z})$ is concentrated on $\mathcal{M}$, we have:
$q(f(\mathbf{z}))>q(\mathbf{z})\quad\Rightarrow\quad\log q(f(\mathbf{z}))>\log q(\mathbf{z}).$
We now Taylor-expand $\log q$ at $\mathbf{z}$ in direction $\mathbf{v}(\mathbf{z})$:
$\log q(f(\mathbf{z}))=\log q(\mathbf{z})+\left\langle\nabla\log q(\mathbf{z}),\mathbf{v}(\mathbf{z})\right\rangle+R,$
where $R=\frac{1}{2}\mathbf{v}^{\top}(\mathbf{z})\nabla^{2}\log q(\xi)\mathbf{v}(\mathbf{z})$, for some $\xi$ between $\mathbf{z}$ and $f(\mathbf{z})$. Since $f$ is contractive, $\|\mathbf{v}(\mathbf{z})\|$ is small and $R=o(\|\mathbf{v}(\mathbf{z})\|)$. Therefore:
$\left\langle\nabla\log q(\mathbf{z}),\mathbf{v}(\mathbf{z})\right\rangle>-R\quad\Rightarrow\quad\left\langle\nabla\log q(\mathbf{z}),\mathbf{v}(\mathbf{z})\right\rangle>0.$
Theorem A.6 (Convergence to latent-space modes).
Let Assumptions 1–4 hold. Let $\mathbf{z}_{0}\in\Omega$, and define the iterative sequence:
$\mathbf{z}_{t+1}:=f(\mathbf{z}_{t})=E(D(\mathbf{z}_{t}))\quad\text{for all }t\geq 0.$
Then:
  1. 1.
    The sequence $\{\mathbf{z}_{t}\}$ converges exponentially fast to a unique fixed point $\mathbf{z}^{*}\in\mathcal{M}$, satisfying $f(\mathbf{z}^{*})=\mathbf{z}^{*}$.
  2. 2.
    At the limit point $\mathbf{z}^{*}$, we have $\nabla\log q(\mathbf{z}^{*})=0$.
  3. 3.
    Moreover, $\mathbf{z}^{*}$ is a local maximum of the density $q$: the Hessian satisfies
    $\nabla^{2}\log q(\mathbf{z}^{*})\prec 0,$
    meaning that the Hessian is negative definite.
Proof.
[…]
  • There exists a unique fixed point $\mathbf{z}^{*}\in\Omega$ such that $f(\mathbf{z}^{*})=\mathbf{z}^{*}$.
  • For any initial $\mathbf{z}_{0}\in\Omega$, the iterates satisfy:
    $\|\mathbf{z}_{t}-\mathbf{z}^{*}\|\leq L^{t}\|\mathbf{z}_{0}-\mathbf{z}^{*}\|\to 0\quad\text{as }t\to\infty.$
Step 2: Stationarity of the limit. At the fixed point $\mathbf{z}^{*}$, we have:
$\mathbf{v}(\mathbf{z}^{*})=f(\mathbf{z}^{*})-\mathbf{z}^{*}=\mathbf{0}.$
Assume for contradiction that $\nabla\log q(\mathbf{z}^{*})\neq\mathbf{0}$. By continuity of $\nabla\log q$, there exists a neighborhood $U\ni\mathbf{z}^{*}$ where $\nabla\log q(\mathbf{z})$ stays close to $\nabla\log q(\mathbf{z}^{*})$. Then for nearby $\mathbf{z}$, Lemma A.5 implies:
$\left\langle\nabla\log q(\mathbf{z}),\mathbf{v}(\mathbf{z})\right\rangle>0.$
But continuity of $f$ and the residual $\mathbf{v}(\mathbf{z})\to\mathbf{0}$ implies the ascent direction vanishes at $\mathbf{z}^{*}$, which contradicts the assumption that $\nabla\log q(\mathbf{z}^{*})\neq\mathbf{0}$. Hence, we conclude:
$\nabla\log q(\mathbf{z}^{*})=\mathbf{0}.$
Step 3: Local maximality. By Lemma A.5 and Assumption 4, the sequence $q(\mathbf{z}_{t})$ is strictly increasing and converges to $q(\mathbf{z}^{*})$. If $\mathbf{z}^{*}$ were a saddle point or local minimum, there would exist a direction $\mathbf{u}\in\mathbb{R}^{d}$ such that the second-order Taylor expansion yields:
$\left.\frac{d^{2}}{dt^{2}}\log q(\mathbf{z}^{*}+t\mathbf{u})\right|_{t=0}=\mathbf{u}^{\top}\nabla^{2}\log q(\mathbf{z}^{*})\mathbf{u}\geq 0,$
which contradicts the fact that $\mathbf{z}_{t}\to\mathbf{z}^{*}$ through ascent directions. Therefore, all second directional derivatives must be negative, implying:
$\nabla^{2}\log q(\mathbf{z}^{*})\prec 0.$
That is, $\mathbf{z}^{*}$ is a strict local maximum of $q$. ∎
A.2 Proposition 1
Proposition A.7.
We report here a detailed formulation of Theorem 1 alongside its proof.
Let $f=E\circ D:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the composition of an autoencoder’s decoder $D$ and encoder $E$, and define the residual vector field $\mathbf{v}(\mathbf{z}):=f(\mathbf{z})-\mathbf{z}$. Consider the reconstruction loss
$L(\mathbf{z}):=\|f(\mathbf{z})-\mathbf{z}\|^{2}.$
Then, the iteration $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ corresponds locally to gradient descent on $L$ (i.e., $\mathbf{v}(\mathbf{z})\propto-\nabla_{\mathbf{z}}L(\mathbf{z})$) if either of the following conditions holds:
  1. 1.
    The Jacobian $J_{f}(\mathbf{z})\approx\mathrm{Id}$ (i.e., $f$ is locally an isometry).
  2. 2.
    $\mathbf{z}$ is near an attractor point $\mathbf{z}^{*}$ such that $f(\mathbf{z}^{*})=\mathbf{z}^{*}$ and $J_{f}(\mathbf{z}^{*})\approx 0$.
Proof.
We begin by expanding the loss:
$L(\mathbf{z})=\|f(\mathbf{z})-\mathbf{z}\|^{2}=\langle f(\mathbf{z})-\mathbf{z},f(\mathbf{z})-\mathbf{z}\rangle.$
Let $\mathbf{v}(\mathbf{z}):=f(\mathbf{z})-\mathbf{z}$. Then,
$L(\mathbf{z})=\|\mathbf{v}(\mathbf{z})\|^{2}.$
Now compute the gradient of $L$:
$\nabla_{\mathbf{z}}L(\mathbf{z})$ $=\nabla_{\mathbf{z}}\left(\mathbf{v}(\mathbf{z})^{\top}\mathbf{v}(\mathbf{z})\right)$
$=2J_{\mathbf{v}}(\mathbf{z})^{\top}\mathbf{v}(\mathbf{z}),$
where $J_{\mathbf{v}}(\mathbf{z})=J_{f}(\mathbf{z})-\mathrm{Id}$ is the Jacobian of the residual field.
So we obtain:
$\nabla_{\mathbf{z}}L(\mathbf{z})=2(J_{f}(\mathbf{z})-\mathrm{Id})^{\top}\mathbf{v}(\mathbf{z}).$
Therefore, if we ask whether $\mathbf{v}(\mathbf{z})\propto-\nabla_{\mathbf{z}}L(\mathbf{z})$, we need:
$\mathbf{v}(\mathbf{z})\propto-(J_{f}(\mathbf{z})-\mathrm{Id})^{\top}\mathbf{v}(\mathbf{z}),$
which simplifies to:
$[(J_{f}(\mathbf{z})-\mathrm{Id})^{\top}+\alpha\mathrm{Id}]\mathbf{v}(\mathbf{z})=0$
for some $\alpha>0$. This holds approximately in the two special cases:
  • Isometry: If $J_{f}(\mathbf{z})\approx\mathrm{Id}$, then $\nabla L(\mathbf{z})\approx 0$, so $\mathbf{v}(\mathbf{z})$ is nearly stationary—i.e., we are at or near a local minimum.
  • Attractor: If $\mathbf{z}\approx\mathbf{z}^{*}$ with $f(\mathbf{z}^{*})=\mathbf{z}^{*}$ and $J_{f}(\mathbf{z}^{*})\approx 0$, then:
    $\nabla L(\mathbf{z}^{*})\approx-2\mathbf{v}(\mathbf{z}^{*})=0,$
    and thus $\mathbf{z}^{*}$ is a fixed point and local minimum of the loss.
In these two cases cases, the dynamics of $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ follow the direction of steepest descent of $L$, up to scaling. Higher order terms dominate influence the dynamics in the general case.
A.3 Proposition 3.2
We report here a detailed formulation of Theorem 3.2 alongside its proof.
Proposition A.8.
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a differentiable function with a fixed point $\mathbf{z}^{*}\in\mathbb{R}^{n}$ such that $f(\mathbf{z}^{*})=\mathbf{z}^{*}$. Assume the Jacobian $J_{f}(\mathbf{z}^{*})$ has spectral norm $\|J_{f}(\mathbf{z}^{*})\|_{\sigma}=\rho<1$. Then, for initial point $\mathbf{z}_{0}$ sufficiently close to $\mathbf{z}^{*}$, the sequence $\{\mathbf{z}_{t}\}$ defined by the iteration $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ satisfies:
$\|\mathbf{z}_{t}-\mathbf{z}^{*}\|\leq\|\mathbf{e}_{0}\|\cdot\rho^{t},$
and the number of iterations $T$ needed to reach an error $\|\mathbf{z}_{T}-\mathbf{z}^{*}\|\leq\epsilon$ is bounded by:
$T\geq\frac{\log(\epsilon/\|\mathbf{e}_{0}\|)}{\log(\rho)}.$
Proof.
Let $\mathbf{e}_{t}=\mathbf{z}_{t}-\mathbf{z}^{*}$ denote the error at iteration $t$. We are given that $f$ is differentiable around $\mathbf{z}^{*}$, and that $f(\mathbf{z}^{*})=\mathbf{z}^{*}$. Applying the first-order Taylor expansion of $f$ around $\mathbf{z}^{*}$, we obtain:
$f(\mathbf{z}_{t})=f(\mathbf{z}^{*})+J_{f}(\mathbf{z}^{*})(\mathbf{z}_{t}-\mathbf{z}^{*})+R(\mathbf{z}_{t}),$
where $R(\mathbf{z}_{t})$ is the Taylor remainder satisfying $\|R(\mathbf{z}_{t})\|=o(\|\mathbf{z}_{t}-\mathbf{z}^{*}\|)$. Since $f(\mathbf{z}^{*})=\mathbf{z}^{*}$, this becomes:
$\mathbf{z}_{t+1}=f(\mathbf{z}_{t})=\mathbf{z}^{*}+J_{f}(\mathbf{z}^{*})(\mathbf{z}_{t}-\mathbf{z}^{*})+R(\mathbf{z}_{t}).$
Thus, the error evolves as:
$\mathbf{e}_{t+1}=\mathbf{z}_{t+1}-\mathbf{z}^{*}=J_{f}(\mathbf{z}^{*})\mathbf{e}_{t}+R(\mathbf{z}_{t}).$
For $\mathbf{z}_{t}$ sufficiently close to $\mathbf{z}^{*}$, we can neglect the higher-order term $R(\mathbf{z}_{t})$ in comparison to the leading linear term. Hence, the error evolves approximately as:
$\|\mathbf{e}_{t+1}\|\leq\|J_{f}(\mathbf{z}^{*})\|_{\sigma}\cdot\|\mathbf{e}_{t}\|+o(\|\mathbf{e}_{t}\|).$
By continuity of $f$, there exists a neighborhood $\mathcal{U}$ of $\mathbf{z}^{*}$ where $\|J_{f}(\mathbf{z})\|_{\sigma}\leq\rho+\delta<1$ for some $\delta>0$. Choosing $\mathbf{z}_{0}\in\mathcal{U}$ and letting $R(\mathbf{z}_{t})=0$ (first-order approximation), the error satisfies:
$\|\mathbf{e}_{t}\|\leq\|\mathbf{e}_{0}\|\cdot\rho^{t}.$
We now ask: for which $T$ does $\|\mathbf{e}_{T}\|\leq\epsilon$? We solve:
$\|\mathbf{e}_{0}\|\cdot\rho^{T}\leq\epsilon.$
Dividing both sides by $\|\mathbf{e}_{0}\|$ and taking logarithms yields:
$\rho^{T}\leq\frac{\epsilon}{\|\mathbf{e}_{0}\|}\quad\Rightarrow\quad T\cdot\log(\rho)\leq\log\left(\frac{\epsilon}{\|\mathbf{e}_{0}\|}\right).$
Since $\log(\rho)<0$, we reverse the inequality when dividing:
$T\geq\frac{\log(\epsilon/\|\mathbf{e}_{0}\|)}{\log(\rho)}.$
This proves that the error converges exponentially and that the number of iterations required to reach error $\epsilon$ is lower bounded by the expression above. ∎
Appendix B Additional experiments
B.1 Bias at initialization
As stated in the method section of the main paper, To empirically measure to what extent model are contractive at initialization we sample $N=1000$ samples from $\mathcal{N}(0,I)$, and we map them trough 12 different vision backbones, initialized randomly across 10 seeds. We measure the ratio between the input variance $\sigma_{in}=1$ and output variance $\sigma_{out}$ and we report in Figure 6 .
Figure 6: Contraction at initialization Variance preserving ratio at initialization of torchvision models: all models considered have a ratio $<1$, indicating that the map at initialization is contractive.
B.2 Qualitative results: data-free weight probing
We report additional qualitative results from the data free weight probing experiment in section 4.1, in Figures 7,8,9, observing that the reconstructions from attractors as a functions of the number of atoms used, are superior with respect to the orthogonal basis baseline.
(a) 256 atoms
(b) 1024 atoms
(c) 2048 atoms
Figure 7: Visualization of reconstructions from the data-free sample recovery experiment of Figure 4: Visualizing on Laion2B reconstructions of five random samples. First row: input samples; second row: reconstructions from orthogonal basis; third row reconstructions from attractors of noise.
(a) 256 atoms
(b) 1024 atoms
(c) 2048 atoms
Figure 8: Visualization of reconstructions from the data-free sample recovery experiment of Figure 4: Visualizing on Imagenet1k reconstructions of five random samples. First row: input samples; second row: reconstructions from orthogonal basis; third row reconstructions from attractors of noise.
(a) 256 atoms
(b) 1024 atoms
(c) 2048 atoms
Figure 9: Visualization of reconstructions from the data-free sample recovery experiment of Figure 4: Visualizing on Camelyon17 reconstructions of five random samples. First row: input samples; second row: reconstructions from orthogonal basis; third row reconstructions from attractors of noise.
B.3 Histograms OOD detection
We report in Figure 10 additional histograms of In-Distribution vs Out-Of-Distribution scores, corresponding to the experiment in Section 4.2.
Figure 10: Histograms from OOD detections
B.4 Analysis speed of convergence
In figure 11 we report an empirical validation of proposition 3.2, for a 2 dimensional bottleneck convolution autoencoder trained on MNIST, by sampling points in the training set and measuring their convergence speed in terms of number of iteration. We remark that assumptions for the proposition don’t hold for this model, as the farer an initial condition is from an attractor the more higher order terms will dominate in the dynamics, making the evaluation of the spectrum of the Jacobian at the attractor insufficient. We observe nevertheless, that the convergence speed bounds still correlates, providing a looser estimate for the number of iteration needed to converge towards an attractors with error $\epsilon$.
Figure 11: Converge speed analysis We plot the convergence speed in terms of number of iterations vs the error (distance to the attractors) of the initial condition on a convolutional autoencoders trained on MNIST. The estimated convergence has a Pearson correlation coefficient of 0.67 to the true number of iterations to converge. Higher error in the estimates occur for farer initial conditions.
B.5 Rank of attractor matrix
In Figure 12 we report an alternative measure of generalization of the attractors computed in the experiment in Figure 2 and in Figure 3. Specifically, we consider the matrix $\mathbf{X}^{*}=D(\mathbf{Z})^{*}$, and we compute its singular value decomposition $\mathbf{U}^{*}diag(\mathbf{s}^{*})\mathbf{V}^{*}=\mathbf{X}^{*}$. We define the generalization entropy as the number of eigenvectors needed to explain $90\%$ of the variance of the matrix. We report this measure for both the models employed in the experiment in Figure2 and in Figure 3, showing that attractors are more expressive as the model generalized better, as well during training.
Figure 12: Generalization increase ranks of attractors matrix: entropy rank as a function of the bottleneck dimension (left) and during training, as a function of the number of epochs (right). Trasnitiong from memorization to generalization
B.6 Memorization overfitting regime
In this section, we test the memorization capabilities of models in different (strong) overparametrization regimes, similarly as tested in [24] for diffusion models and in [42, 56] for the extremely overparametrized case (network trained on few samples). We remark that this case corresponds to an overfitting regime of the network, as opposed as underfitting (over-regularized) regime showed in the main paper. We train a convolutional autoencoder on subsamples of the CIFAR,MNIST,FashionMNIST datasets, with bottleneck dimension $k=128$ and weight decay $41e-$ 4. In Figure 13 we plot the memorization coefficient as a function of the dataset size, showing that networks trained on less data( strong overparametrization) tend to memorize more the training data.
Figure 13: Strong overparametrization favors memorization: Memorization coefficient as a function of the dataset size on CIFAR,MNIST,FashionMNIST, showing that networks trained on less data (strongly overparametrized) tend to memorize more the training data.
B.7 Data-free weight probing: additional results
We replicate the experiment in 4.1, on the larger XL version of the Stable diffusion architecture [46], observing that attractors from noise still form an informative dictionary of signals which scales better w.r.t, to a random orthogonal basis.
Figure 14: Data-free weight probing Stable diffusion XL AE : the results on the larger version of the table diffusion autoencoder confirm that attractors from noise form an informative dictionary of signals to reconstruct different datasets.
Table 1: Architecture of the AutoencoderKL in Stable Diffusion
[…]
Encoder
Input $3\times 256\times 256$ image $3\times 256\times 256$
Conv2D $3\to 128$, $3\times 3$, stride 1 $128\times 256\times 256$
ResNet Block $128$ channels $128\times 256\times 256$
Conv2D $128\to 256$, $3\times 3$, stride 2 $256\times 128\times 128$
ResNet Block $256$ channels $256\times 128\times 128$
Conv2D $256\to 512$, $3\times 3$, stride 2 $512\times 64\times 64$
ResNet Block $512$ channels $512\times 64\times 64$
Conv2D $512\to 512$, $3\times 3$, stride 2 $512\times 32\times 32$
ResNet Block $512$ channels $512\times 32\times 32$
Conv2D $512\to 512$, $3\times 3$, stride 2 $512\times 16\times 16$
ResNet Block $512$ channels $512\times 16\times 16$
Conv2D $512\to 512$, $3\times 3$, stride 2 $512\times 8\times 8$
ResNet Block $512$ channels $512\times 8\times 8$
Mean and Log Variance $512\to 4$, $1\times 1$ $4\times 8\times 8$
Sampling Reparameterization trick $4\times 8\times 8$
Decoder
Conv2D $4\to 512$, $3\times 3$, stride 1 $512\times 8\times 8$
ResNet Block $512$ channels $512\times 8\times 8$
Upsample Scale factor 2 $512\times 16\times 16$
ResNet Block $512$ channels $512\times 16\times 16$
Upsample Scale factor 2 $512\times 32\times 32$
ResNet Block $512$ channels $512\times 32\times 32$
Upsample Scale factor 2 $512\times 64\times 64$
ResNet Block $512$ channels $512\times 64\times 64$
Upsample Scale factor 2 $256\times 128\times 128$
ResNet Block $256$ channels $256\times 128\times 128$
Upsample Scale factor 2 $128\times 256\times 256$
ResNet Block $128$ channels $128\times 256\times 256$
Conv2D $128\to 3$, $3\times 3$, stride 1 $3\times 256\times 256$
Output: Reconstructed Image
Appendix C Regularized autoencoder enforce contractiveness
In Table C, we list many AE variants including denoising AEs (DAEs) [52], sparse AEs (SAEs) [37], variational AEs (VAEs)  [26] and other variants [44, 2, 12] and show how their objectives enforce local contractive solutions around training points.
[…]
Standard AE (rank $\leq k$) – (bottleneck dimension $k$) No explicit regularization, but dimensionality constraint induces compression Bottleneck limits Jacobian rank: $\mathrm{rank}(J_{f_{\theta}}(\mathbf{x}))\leq k$
Plain AE with Weight Decay $\|\mathbf{W}_{E}\|_{F}^{2}+\|\mathbf{W}_{D}\|_{F}^{2}$ L2 penalty on encoder and decoder weights Reduces $\|J_{f_{\theta}}(\mathbf{x})\|\leq L_{\sigma}^{2}\|\mathbf{W}_{D}\|\|\mathbf{W}_{E}\|$ by shrinking weight norms
Deep AE with Weight Decay $\sum_{\ell}\|\mathbf{W}_{E}^{(\ell)}\|_{F}^{2}+\|\mathbf{W}_{D}^{(\ell)}\|_{F}^{2}$ Weight decay across multiple layers of deep encoder/decoder Layerwise shrinkage enforces smoother, more contractive composition $J_{f_{\theta}}(\mathbf{x})$
R-Contractive AE [1] $\|J_{f_{\theta}}(\mathbf{x})\|_{F}^{2}$ Penalizes Jacobian of the full map $f_{\theta}=D_{\theta}\circ E_{\theta}$ Encourages stability of reconstructions: $f_{\theta}(\mathbf{x})\approx f_{\theta}(\mathbf{x}+\delta)$
Contractive AE [44] $\|J_{E_{\theta}}(\mathbf{x})\|_{F}^{2}=\sum_{i=1}^{k}\|J_{E_{i}}(\mathbf{x})\|^{2}$ Penalizes encoder Jacobian norm Explicitly enforces local flatness of encoder
Sparse AE (KL, sigmoid) [37] $\sum_{i=1}^{k}\mathrm{KL}(\rho\,\|\,\hat{\rho}_{i}),\quad\hat{\rho}_{i}=\mathbb{E}_{\mathbf{x}}[E_{i}(\mathbf{x})]$ Enforces low average activation under sigmoid Saturated units $\Rightarrow J_{E_{i}}(\mathbf{x})\approx 0\Rightarrow\|J_{E_{\theta}}(\mathbf{x})\|$ small
Sparse AE (L1, ReLU) [37] $\|E_{\theta}(\mathbf{x})\|_{1}$ Promotes sparsity of ReLU activations Inactive ReLU units have zero derivatives $\Rightarrow$ sparse Jacobian $J_{E_{\theta}}(\mathbf{x})$
Denoising AE [52, 2] $\mathbb{E}_{\tilde{\mathbf{x}}\sim q(\tilde{\mathbf{x}}|\mathbf{x})}\|\mathbf{x}-f_{\theta}(\tilde{\mathbf{x}})\|^{2}$ Reconstructs from noisy input For small noise: $f_{\theta}(\mathbf{x})-\mathbf{x}\approx\sigma^{2}\nabla\log p(\mathbf{x})$, a contractive vector field $J_{f_{\theta}}(\mathbf{x})$
Variational AE (VAE)[26] $\mathrm{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z}))$, where $q(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mu(\mathbf{x}),\Sigma(\mathbf{x}))$ Encourages latent distribution to match prior Smooths $\mu(\mathbf{x}),\Sigma(\mathbf{x})$; penalizes sharp variations in encoder $J_{\mu}(\mathbf{x}),J_{\Sigma}(\mathbf{x})$
Masked AE (MAE)[15] $\mathbb{E}_{M}\left[\|M\odot(\mathbf{x}-f_{\theta}(M\odot\mathbf{x}))\|^{2}\right]$ Learns to reconstruct from partial input Promotes invariance to missing entries $\Rightarrow\|J_{f_{\theta}}(\mathbf{x})\|$ low
Table 2: Unified formulation of autoencoder variants as minimizing reconstruction error plus a regularizer that encourages local contractiveness.
Appendix D Additional implementation details
Table 3: Architecture of the autoencoder employed in experiments on CIFAR, MNIST, FashionMNIST datasets
[…]
Encoder
Input $C\times H\times W$ image $C\times H\times W$
Conv2D $C\to d$, $3\times 3$, stride 2, pad 1 $d\times\frac{H}{2}\times\frac{W}{2}$
Conv2D $d\to 2d$, $3\times 3$, stride 2, pad 1 $2d\times\frac{H}{4}\times\frac{W}{4}$
Conv2D $2d\to 4d$, $3\times 3$, stride 2, pad 1 $4d\times\frac{H}{8}\times\frac{W}{8}$
Conv2D $4d\to 8d$, $3\times 3$, stride 2, pad 1 $8d\times\frac{H}{16}\times\frac{W}{16}$
Flatten $8d\cdot\frac{H}{16}\cdot\frac{W}{16}$
Bottleneck Linear layer or projection $z$ (latent code)
Decoder
Unflatten $z\to 8d\times\frac{H}{16}\times\frac{W}{16}$ $8d\times\frac{H}{16}\times\frac{W}{16}$
ConvTranspose2D $8d\to 4d$, $3\times 3$, stride 2, pad 1, output pad 1 $4d\times\frac{H}{8}\times\frac{W}{8}$
ConvTranspose2D $4d\to 2d$, $3\times 3$, stride 2, pad 1, output pad 1 $2d\times\frac{H}{4}\times\frac{W}{4}$
ConvTranspose2D $2d\to d$, $3\times 3$, stride 2, pad 1, output pad 1 $d\times\frac{H}{2}\times\frac{W}{2}$
ConvTranspose2D $d\to C$, $3\times 3$, stride 2, pad 1, output pad 1 $C\times H\times W$
Output: Reconstructed Image
For the experiments in Figures 2,3, 13 we considered convolutional autoencoders with architecture specified in Table D. We train the models with Adam [25] with learning rate $45e-$ 4, linear step learning rate scheduler, and weight decay $41e-$ 4 for $500$ epochs. For the experiements in section 4.1, we use the pretrained autoencoder of [46]. We report in Table LABEL:tab:arch_autoencoderKL the architectural details and we refer to the paper for training details. For the experiments in section 4.2, we use the pretrained autoencoder of [15], considering the base model. We refer to the paper for training details. For computing attractors in experiments in Figures 2,3, 13, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-6$ or reaching $t=3000$. Similarly in experiments in Sections 4.1, 4.2, we compute the iterations $\mathbf{z}_{t+1}=f(\mathbf{z}_{t})$ until $\|f(\mathbf{z}_{t+1})-f(\mathbf{z}_{t})\|_{2}^{2}<1e-5$ or reaching $t=500$. All experiments are performed on a GPU NVIDIA 3080TI in Python code, using the PyTorch library [38].